OpenStudy (anonymous):

use a maclaurin series representation of the integrand from 0 to 1 of cos(x^2) to approximate the integral with an error of no more than .0001

OpenStudy (anonymous):

\[\cos x=1-\frac{ x ^{2} }{ 2! }+\frac{ x ^{4} }{ 4! }-\frac{ x ^{6} }{ 6! }+ ...\] put in x^2 in place of x and integrate term bu term until the integral is less than the desired accuracy

OpenStudy (anonymous):

hint: you'll only need to do 2 terms

OpenStudy (anonymous):

ok im still confused though

OpenStudy (anonymous):

about...

OpenStudy (anonymous):

where do i go after integrating

OpenStudy (anonymous):

they want you to approximate \[\int\limits\limits_{0}^{1}\cos x ^{2}dx\] to the nearest .0001 by switching out the integrand with its MacLaurin series represetation. so we get \[\int\limits\limits_{0}^{1}\cos x ^{2}dx=\int\limits\limits_{0}^{1}\left( 1-\frac{ x ^{4} }{ 2! }+\frac{ x ^{8} }{ 4! }- ... \right) dx\] so integrate term by term until the numbers in the 4th decimal place stop changing

OpenStudy (anonymous):

ok thank you for your help